# Physics-Driven Local-Whole Elastic Deformation Modeling for Point Cloud Representation Learning

### v2

@@ Abstract
Existing point cloud representation learning methods primarily rely on data-driven strategies to extract geometric information from large amounts of scattered data. However, most methods focus solely on the spatial distribution features of point clouds while overlooking the relationship between local information and the whole structure, which limits the accuracy of point cloud representation. Local information reflect the fine-grained variations of an object, while the whole structure is determined by the interaction and combination of these local features, collectively defining the object’s shape. In real-world, objects undergo deformation under external forces, and this deformation gradually affects the whole structure through the propagation of forces from local regions, thereby altering the object’s geometric features. Therefore, the appropriate introduction of physics-driven mechanism can effectively compensate for the limitations of data-driven methods in structural modeling and significantly enhance the generalization and interpretability of point cloud representations in downstream tasks such as understanding and recognition. Inspired by this, we incorporate a physics-driven mechanism into the data-driven method to learn fine-grained features in point clouds and model the structural relationship between local regions and the whole shape. Specifically, we design a dual-task encoder-decoder framework that combines the geometric modeling capability of data-driven implicit fields with physics-driven elastic deformation. Through the integration of physics-based loss functions, the framework is guided to predict localized deformation and explicitly capture the correspondence between local structural changes and whole shape variations. Experimental results show that our method outperforms existing approaches in object classification and segmentation, demonstrating its effectiveness.
Index Terms:
Point Cloud Representation, Physics-driven, Implicit Field, Self-supervised Learning.
@@ I Introduction
Figure 1: (a) Visualization of deformation and force distribution under an external force $f$ for Lamp A and B. Colors represent force magnitude, with red circles indicating regions of significant deformation. (b) Clustering results with and without Physics-Informed (PI) modeling.
Existing methods primarily rely on data-driven approaches to capture geometric distributions of point clouds. In supervised point cloud representation learning, typical network architectures, such as PointNet [5], use shared MLPs to process unordered point sets and directly learn the feature representations of each point in the point cloud to capture the whole geometric distribution. PointNet++ [6] builds upon this by introducing a hierarchical structure to capture local features, allowing for better capture of structural information in local regions. DGCNN [7] utilizes dynamic graph convolutional networks to capture local geometric structural features. In unsupervised point cloud representation learning, generative methods [8, 9, 10, 11] learn latent representations of the data by reconstructing the geometric structure of the point cloud using generative models such as autoencoder (AE) [12], generative adversarial networks (GAN) [13], or autoregressive (AR) [10]. Discriminative methods [14, 15, 16, 17] typically construct tasks by comparing feature differences between different augmented samples or categories, focusing more on learning the whole distribution patterns [18, 19] of point clouds in the feature space. Despite the significant progress made by these methods in point cloud representation learning, they often overlook the interaction between local information and whole structure, failing to explicitly model their interrelationship.
@@
In nature, the shape of objects is often governed by the forces acting upon them. When an object is subjected to external forces, slight deformations first occur in its local regions. These deformations propagate through structural connections, causing global shape changes. This transition from “local stress” to “global response” reveals the inherent coupling relationships among the parts of an object. Meanwhile, recent research [20] has shown that physical modeling not only provides interpretability and inductive capability, but can also be embedded into deep networks as structural constraints to enhance the performance of data-driven methods on downstream detection tasks. As shown in Figure 1, incorporating the physics-driven approach enables more effective discrimination of fine-grained features both intra-class and inter-class. Inspired by this, we incorporate a physics-driven mechanism into the data-driven method, which captures the deformation effects between local and whole shapes by introducing the force propagation mechanism from elasticity theory. Specifically, we design a dual-task encoder-decoder framework. The encoder is responsible for learning the geometric information and physical properties of object from both the point cloud data and its corresponding tetrahedral mesh representation. The learned features are then input into two decoders: one decoder models the whole geometric shape of the point cloud using implicit fields, while the other decoder uses data-fidelity loss function and physics-informed loss function to predict local deformations, simulating the force propagation relationship between local and whole structures. The main contributions of our method are as follows:
$\bullet$ We propose a novel self-supervised learning method that explicitly captures the relationship between local features and whole structure in point clouds by combining physical laws with geometric modeling. This is also the first self-supervised learning method that leverages the mechanism of force propagation to model the relationship between local and whole structures of objects.
$\bullet$ We designed two key loss functions to learn the physical characteristics of objects: a data-fidelity loss and a physics-informed loss. These two loss functions work synergistically during the optimization process to enforce physical consistency in the point cloud.
@@ II-B Unsupervised Point Cloud Representation Learning
Nowadays, numerous unsupervised learning methods have emerged. Generative methods learn features through self-reconstruction [26, 8, 27, 28, 11, 29]. For example, HC-VAE [27] introduces a hierarchical consistency mechanism to progressively generate high-quality point clouds. FoldingNet [28] reconstructs arbitrary point clouds using a graph-based encoder and a folding-based decoder. Recently, recovering missing parts from masked inputs [9, 10, 8, 30] has also become a prominent research focus. MaskSurf [9] designs a transformer-based encoder-decoder network to estimate the position of the masking point and its corresponding normal vector. Point-DAE [8] reconstructs various corrupted point clouds by decomposing the reconstruction of a complete point cloud into detailed local patches and a coarse global shape, thereby alleviating the issue of positional leakage in the reconstruction process.
@@
Discriminative methods guide the model to learn meaningful feature representations by designing tasks that focus on the feature differences between samples. Some methods employ pretext tasks to learn point cloud representations [31, 32, 14, 33, 34, 35]. Jigsaw3D [31] uses 3D puzzles as a self-supervised learning task, allowing the model to learn how to correctly assemble object parts. Poursaeed et al. [36] define the estimation of point cloud rotation angles as a pretext task. Recently, contrastive learning methods [15, 37, 14] have been widely used in 3D self-supervised learning, achieving state-of-the-art performance. PointContrast [14] performs point-level invariant mapping on two different directional views of the same point cloud. It is the first unified framework to study the contrastive paradigm for 3D representation learning. STRL [15] is an extension of the method [38] on 3D point clouds to learn rich representational information by updating between online and target networks. PointSmile [17] guides network learning by maximizing the mutual information between two point clouds. In addition, some researchers have focused on incorporating cross-model [39, 40, 41, 16, 42, 43, 44] methods into discriminative methods. However, these methods often overlook the influence of local structures on whole shape responses and typically rely on large numbers of positive-negative sample pairs and complex strategies, resulting in a learning process that lacks physical interpretability.
@@
Figure 2: The overall architecture of the proposed method is illustrated. It consists of two modules: the implicit feature learning module, which represents the point cloud as a continuous surface, helping the network better learn the whole shape features of the object; and the physics information awareness module, which learns the relationship between local features and whole geometry through the elastic deformation of the object, enhancing the model’s ability to learn local details.
In this work, we incorporate a physics-driven mechanism into the data-driven method that explicitly models deformation propagation from local forces to whole structures, enabling physically consistent and interpretable point cloud understanding. This also leads to more robust and generalizable features for downstream tasks.
@@ III Method
III-A Implicit Feature Learning
For a point cloud dataset $\mathcal{D}=\left\{{{P}_{i}}\right\}_{i=1}^{\left|\mathcal{D}\right|}$, we use encoder ${\Phi}_{\theta}$} and decoder ${\Psi}_{\eta}$} to learn the implicit features of the point cloud ${P_{i}}=\left\{{p_{i}^{j}\in{\mathbb{R}^{3}}|j=1,2,...,n}\right\}$. The encoder ${\Phi}_{\theta}$} takes the $P_{i}$} as input and maps it to an $m$-dimensional latent vecoter $z_{\eta}$}, represented as ${{\Phi}_{\theta}}:{{\mathbb{R}}^{n\times 3}}\to{{\mathbb{R}}^{m}},m\ll n$, where we use the network backbones of PointNet [5] or DGCNN [7] as ${\Phi}_{\theta}$} to learn the point cloud features, respectively. Subsequently, we concatenate latent vector $z_{\eta}$} with an arbitrary query point $q$ in 3D space to obtain feature vector $z_{\eta}^{q}={z_{\eta}}\oplus q$. Then, the decoder ${\Psi}_{\eta}$} takes $z_{\eta}^{q}$ as input and output an implicit function $\mathbb{S}={{\Psi}_{\eta}}\left(q,{{\Phi}_{\theta}}\left({{P}_{i}}\right)\right)=\left\{{\delta_{i}}|{\delta_{i}}\left({z_{\eta}^{q}}\right)\in\mathbb{R},\forall q\in{{\mathbb{R}}^{3}}\right\}$. ${\delta_{i}}\left({z_{\eta}^{q}}\right)$ represents the shortest distance from the query point $q$ to the point cloud $P_{i}$}. The decoder ${\Psi}_{\eta}$} consists of 3-layer of MLP and outputs a 1-dimensional feature vector.
@@ III-A Implicit Feature Learning
III-B Physical Information Awareness
@@
For a point cloud ${P}_{i}$}, we first use 3D Delaunay [46] to convert the ${P}_{i}$} into an object composed of multiple tetrahedra. It is an extension of Delaunay triangulation (which generates triangular meshes in 2D) to 3D space, producing a mesh composed of tetrahedral elements. Compared to hexahedral meshes, tetrahedral meshes better fit free-form objects and support complex topological structures, providing the volumetric elements required for physical modeling, such as deformation gradients and stress tensor computation. Meanwhile, directly using point coordinates lacks structural information, making it difficult to establish the strain–stress relationships necessary for constructing the subsequent physics-informed loss ${\mathcal{L}}_{pi}$. The advantages of using tetrahedral structures over point-based representations will be further validated in the ablation studies. The meshing process of the point cloud is illustrated in Figure 3.
@@
Figure 3: The point cloud is first converted into a rough mesh form through Delaunay triangulation. Then, we removed the larger tetrahedral meshes, further refining them into the final tetrahedral representation.
@@
Subsequently, the tetrahedral mesh data is processed by the mesh processor $\mathrm{T}$, which encodes the local geometric and topological structure of each tetrahedron into a unified format suitable for input to the encoder ${\Phi}_{\theta}$}, where the mesh processor $\mathrm{T}$ consists of 3-layer of convolutional networks, serving as a bridge between the mesh data and encoder ${\Phi}_{\theta}$}. The encoder ${\Phi}_{\theta}$} maps the processed data to the $m$-dimensional latent space to obtain the latent vector $z_{phys}$}. $z_{phys}$} is concatenated with the deformation parameters $d$ as the input $z_{phy}^{d}$ to decoder ${\Psi}_{phys}$}, enabling the model to learn local-to-whole deformation patterns under varying physical conditions during training. The $d$ includes standard physical parameters commonly used in static linear elasticity theory, such as Lamé constants $\lambda$ and $\mu$, elastic modulus $E$, Poisson’s ratio $\nu$, and external forces $f$. These parameters have clear physical meanings and can effectively characterize stress responses of typical elastic materials under small deformations. Considering the differences in point cloud density and actual deformation scale, we further apply an adaptive adjustment to the parameters in $d$ based on the size of the generated tetrahedral mesh and the initial deformation magnitude. This reduces reliance on manual configuration and enhances the robustness and generalization ability of the proposed method. Decoder ${\Psi}_{phys}$} is composed of 5-layer of MLP, outputting a feature vector of size $n\times 3$, where $n$ is the number of input points in the point cloud.
@@
Based on the configuration of deformation parameters $d$, we obtain the ground-truth displacement field $u$ using the finite element analysis method (FEM) [47]. Finally, based on the displacement field $\hat{u}$ predicted by the network and the ground-truth displacement field $u$ obtained through computation, we design two loss functions: the data-fidelity loss and the physics-informed loss. The data-fidelity loss guides the network to predict the displacement of each point in the deformed point cloud, explicitly modeling the relationship between local deformations and the whole geometric shape, thereby revealing the influence of local features on the whole structure. The physics-informed loss derives the external force distribution from the predicted local deformation data and compares it with the true external force distribution, ensuring that the network’s output adheres to the physical laws. Below, we provide a detailed description of these two loss functions.
Data-fidelity Loss. Its objective is to guide the network in accurately predicting the displacement field $\hat{u}$ deformed tetrahedral mesh object and comparing it with the ground truth displacement field u{u}, enabling it to precisely predict local geometric changes. In calculating elastic deformation, the finite element method is a common approach to solving elastic dynamics models in contact problems. Given $n$ node positions $x$, a finite element mass matrix $M$, and a hyper-elastic deformation energy $E(x)$, the solution to the contact problem can be obtained using the extremization method of the action:
@@
where $S(x)$ represents the action of the system over the time interval $\left[{0,T}\right]$, describing the total energy of the object. $\frac{1}{2}{{{\dot{x}}}^{T}}M\dot{x}$ denotes the kinetic energy, $E(x)$ represents the potential energy (elastic deformation potential energy) and ${{x}^{T}}\left({{f}_{e}}+{{f}_{d}}\right)$ corresponds to the external force action. ${f}_{e}$} are external forces and ${f}_{d}$} are dissipative frictional forces.
The above describes the general analysis of dynamic problems. To simulation and analyze the deformation of an object under static linear elastic forces within the spatial domain $\Omega$}, we use FEM to establish the stress-strain equation and solve for the displacement field $u$:
@@
where $\Omega$} represents the volume region of the object, and $S(u)$ computes the total energy of the object’s deformation. $\frac{1}{2}\sigma\colon\varepsilon(u)$ represents the energy stored in the tetrahedral mesh due to elastic deformation. After the mesh is subjected to force, the resulting energy is calculated by the stress $\sigma$ and strain $\varepsilon$. $f\cdot u$ represents the potential energy associated with external body forces under the displacement field, which is the work done by the external force under the displacement field $u$.
We determine the force application direction by computing the object’s inertia matrix [48] to ensure that the deformation direction better aligns with the object’s geometric distribution, avoiding potential unreasonable force application caused by manual selection. The eigenvectors of the inertia matrix reflect the primary distribution trends of the object in different directions, providing an optimal force application direction that conforms to its geometric structure, thereby ensuring a more reasonable force application. Finally, we design the following data-fidelity loss:
@@
Physics-informed Loss. Based on the predicted displacements of tetrahedral vertices by the network, we compute the external force residuals by deriving stress tensors from predicted displacements using Hooke’s law [49] and evaluating their divergence based on the equilibrium equation [50]. This local deformation is propagated through the entire mesh via force equilibrium constraints, enforcing consistency across connected tetrahedra and shaping the whole displacement field $\hat{u}$. First, each tetrahedron is composed of four vertices, denoted as ${{x}_{1}},{{x}_{2}},{{x}_{3}},{{x}_{4}}$ (initial coordinates) and ${{{x}^{\prime}}_{1}},{{{x}^{\prime}}_{2}},{{{x}^{\prime}}_{3}},{{{x}^{\prime}}_{4}}$ (coordinates after deformation). We construct the initial shape matrix ${X}=\left[{{x_{2}}-{x_{1}}\quad{\rm{}}{x_{3}}-{x_{1}}\quad{\rm{}}{x_{4}}-{x_{1}}}\right]$ and the shape matrix ${X^{\prime}}=\left[{{{x^{\prime}}_{2}}-{{x^{\prime}}_{1}}\quad{\rm{}}{{x^{\prime}}_{3}}-{{x^{\prime}}_{1}}\quad{\rm{}}{{x^{\prime}}_{4}}-{{x^{\prime}}_{1}}}\right]$ learned by the network. Then, we construct the deformation gradient $F$ based on the shape matrices $X$ and $X^{\prime}$:
@@
where ${F}^{T}$ is the transpose matrix of the deformation gradient, and $I$ is the identity matrix.
In elasticity, stress $\sigma$ and strain $\varepsilon$ are related through constitutive relations. For linear elastic materials, Hooke’s law is usually used for description. In the three-dimensional cases, the stress tensor is calculated as:
@@
where $\lambda=\frac{E\nu}{\left(1+\nu\right)\left(1-2\nu\right)}$ and $\mu=\frac{E}{2\left(1+\nu\right)}$ is the Lamé constant, which is related to the elastic modulus $E$ and Poisson’s ratio $\nu$. $Tr\left(\varepsilon\right)$ is the trace of the strain tensor, representing the volumetric deformation. These parameters, along with the external force $f$ , determine how the object will deform.
In three-dimensional linear elasticity, the equilibrium equation is $\nabla\cdot\sigma+f=0$. Therefore, our physics-informed loss term is:

### v3

@@ Abstract
Existing point cloud representation learning methods primarily rely on data-driven strategies to extract geometric information from large amounts of scattered data. However, most methods focus solely on the spatial distribution features of point clouds while overlooking the relationship between local information and the whole structure, which limits the accuracy of point cloud representation. Local information reflect the fine-grained variations of an object, while the whole structure is determined by the interaction and combination of these local features, collectively defining the object’s shape. In real-world, objects undergo deformation under external forces, and this deformation gradually affects the whole structure through the propagation of forces from local regions, thereby altering the object’s geometric features. Therefore, appropriately introducing a physics-driven mechanism to capture the topological relationships between local parts and the whole object can effectively mitigate for the limitations of data-driven point cloud methods in structural modeling, and enhance the generalization and interpretability of point cloud representations for downstream tasks such as understanding and recognition. Inspired by this, we incorporate a physics-driven mechanism into the data-driven method to learn fine-grained features in point clouds and model the structural relationship between local regions and the whole shape. Specifically, we design a dual-task encoder-decoder framework that combines the geometric modeling capability of data-driven implicit fields with physics-driven elastic deformation. Through the integration of physics-based loss functions, the framework is guided to predict localized deformation and explicitly capture the correspondence between local structural changes and whole shape variations. Experimental results show that our method outperforms existing approaches in object classification and segmentation, demonstrating its effectiveness.
@@ I Introduction
Figure 1: (a) Intra-class deformation and force distribution under an external force $f$ for Lamp A and B (colors indicate force magnitude; highlighted regions show significant deformation). (b) Inter-class clustering visualization in the embedding space across multiple categories, comparing models without and with Physics-Informed (PI) modeling. Note that (a) and (b) correspond to different tasks and the color encodings are not shared.
Existing methods primarily rely on data-driven approaches to capture geometric distributions of point clouds. In supervised point cloud representation learning, typical network architectures, such as PointNet [5], use shared MLPs to process unordered point sets and directly learn the feature representations of each point in the point cloud to capture the whole geometric distribution. Building on this, PointNet++ [6] aggregates local features hierarchically within fixed spatial neighborhoods via layered sampling and grouping, whereas DGCNN [7] dynamically constructs a $k$-NN graph in the feature space and learns with EdgeConv, allowing neighborhood relations to adapt as features are updated. In unsupervised point cloud representation learning, generative methods [8, 9, 10, 11] learn latent representations of the data by reconstructing the geometric structure of the point cloud using generative models such as autoencoder (AE) [12], generative adversarial networks (GAN) [13], or autoregressive (AR) [10]. Discriminative methods [14, 15, 16, 17] typically construct tasks by comparing feature differences between different augmented samples or categories, focusing more on learning the whole distribution patterns [18, 19, 20] of point clouds in the feature space. Despite the significant progress made by these methods in point cloud representation learning, they often overlook the interaction between local information and whole structure, failing to explicitly model their interrelationship.
@@
In nature, the shape of objects is often governed by the forces acting upon them. When an object is subjected to external forces, slight deformations first occur in its local regions. These deformations propagate through structural connections, causing global shape changes. This transition from “local stress” to “global response” reveals the inherent coupling relationships among the parts of an object. Meanwhile, recent research [21] has shown that physical modeling not only provides interpretability and inductive capability, but can also be embedded into deep networks as structural constraints to enhance the performance of data-driven methods on downstream detection tasks. As shown in Figure 1, incorporating the physics-driven approach enables more effective discrimination of fine-grained features both intra-class and inter-class. Inspired by this, we incorporate a physics-driven mechanism into the data-driven method to extract local and whole structural shape relationships from object deformations. Specifically, we design a dual-task encoder-decoder framework. The encoder is responsible for learning the geometric information and physical properties of object from both the point cloud data and its corresponding tetrahedral mesh representation. The learned features are then input into two decoders: one decoder models the whole geometric shape of the point cloud using implicit fields, while the other decoder uses data-fidelity loss function and physics-informed loss function to predict local deformations, learning the force propagation relationship between local and whole structures. The main contributions of our method are as follows:
$\bullet$ We propose a novel self-supervised learning method that combines physical laws with geometric modeling and exploits force propagation to capture the relationships between local features and whole structures in point clouds.
$\bullet$ We designed two key loss functions: a data-fidelity loss that aligns the predicted displacement field for geometric consistency, and a physics-informed loss that enforces force equilibrium for physical consistency. Together, they regularize deformation behavior in point clouds.
@@ II-B Unsupervised Point Cloud Representation Learning
Nowadays, numerous unsupervised learning methods have emerged. Generative methods learn features through self-reconstruction [27, 8, 28, 29, 11, 30]. For example, HC-VAE [28] introduces a hierarchical consistency mechanism to progressively generate high-quality point clouds. FoldingNet [29] reconstructs arbitrary point clouds using a graph-based encoder and a folding-based decoder. Recently, recovering missing parts from masked inputs [9, 10, 8, 31, 32] has also become a prominent research focus. MaskSurf [9] designs a transformer-based encoder-decoder network to estimate the position of the masking point and its corresponding normal vector. Point-DAE [8] reconstructs various corrupted point clouds by decomposing the reconstruction of a complete point cloud into detailed local patches and a coarse global shape, thereby alleviating the issue of positional leakage in the reconstruction process.
@@
Discriminative methods guide the model to learn meaningful feature representations by designing tasks that focus on the feature differences between samples. Some methods employ pretext tasks to learn point cloud representations [33, 34, 14, 35, 36]. Jigsaw3D [33] uses 3D puzzles as a self-supervised learning task, allowing the model to learn how to correctly assemble object parts. Poursaeed et al. [37] define the estimation of point cloud rotation angles as a pretext task. Recently, contrastive learning methods [15, 38, 14, 39] have been widely used in 3D self-supervised learning, achieving state-of-the-art performance. PointContrast [14] performs point-level invariant mapping on two different directional views of the same point cloud. It is the first unified framework to study the contrastive paradigm for 3D representation learning. STRL [15] is an extension of the method [40] on 3D point clouds to learn rich representational information by updating between online and target networks. PointSmile [17] guides network learning by maximizing the mutual information between two point clouds. In addition, some researchers have focused on incorporating cross-model [20, 41, 42, 43, 16, 44, 45, 46] methods into discriminative methods. However, these methods often overlook the influence of local structures on whole shape responses and typically rely on large numbers of positive-negative sample pairs and complex strategies, resulting in a learning process that lacks physical interpretability.
II-C Physics-Informed Learning for Elasticity
Recent physics-informed learning approaches [21] for elasticity often learn high-fidelity surrogate solvers for FEM-discretized fields by embedding physical laws into neural networks through loss terms or constraints. For example, DFEM [47] couples PINNs with Finite Element Method (FEM) [48] discretization and leverages stiffness matrix constraints to improve the stability and accuracy of 3D elasticity solutions. Neural Modes [49], on the other hand, learns an interpretable low-dimensional modal subspace by minimizing the system’s mechanical energy under equilibrium constraints. In the modeling of discrete structural systems, GNN-based methods (e.g., StructGNN-E [50] and Gulakala et al. [51]) further organize FEM-discretized nodes, edges, and boundary conditions into graphs, and use message passing to predict internal forces or von Mises stress fields. In contrast, we use static linear elasticity as a lightweight yet reliable physical prior to regularize point cloud self-supervised learning. Specifically, we combine data-driven learning with physics-based knowledge to capture how local deformations propagate to global responses under force equilibrium, thereby explicitly modeling the structural relationship between local parts and the whole object. This yields physically consistent and interpretable representations, mitigates the structural modeling limitations of purely data-driven methods, and improves robustness and generalization for downstream tasks.
@@ II-C Physics-Informed Learning for Elasticity
Figure 2: The overall architecture of the proposed method is illustrated. It consists of two modules: the implicit feature learning (IFL) module, which represents the point cloud as a continuous surface, helping the network better learn the whole shape features of the object; and the physics information awareness (PIA) module, which learns the relationship between local features and whole geometry through the elastic deformation of the object, enhancing the model’s ability to learn local details.
@@ III Method
III-A Preliminaries
Given a point cloud ${P_{i}}=\left\{{p_{i}^{j}\in{\mathbb{R}^{3}}|j=1,2,...,n}\right\}$ representing a shape, we denote the implicit field query set as $Q=\left\{{{q_{k}}}\right\}_{k=1}^{K}$, where ${q_{k}}\in{\mathbb{R}}{{}^{3}}$. The backbone encoder ${\Phi}_{\theta}$} extracts a global latent code $z$ (dimension $m$), which is consumed by two decoders: an implicit decoder ${\Psi}_{\eta}$} for occupancy prediction and a deformation decoder ${\Psi}_{phys}$} for physics-driven shape response. For the physics branch, we represent the volumetric domain by a tetrahedral mesh $\Omega_{h}$} and use a mesh processor $\mathrm{T}$ to produce tetrahedron features compatible with ${\Phi}_{\theta}$}, yielding a latent code $z_{phys}$}. We then introduce $\tau$ as a physics-driven knowledge signal that defines the deformation configuration and thereby guides the network to learn physically meaningful deformation behavior. Subsequently, we concatenate $\tau$ with $z_{phys}$} to form a conditioned representation (i.e., ${z_{phys}}\oplus d$), which is fed into ${\Psi}_{phys}$} to predict the displacement field and obtain the physics-driven shape response. In our implementation, $\tau$ encodes the key ingredients of linear elasticity, including material properties (e.g., Young’s modulus $E$ and Poisson’s ratio $\nu$, equivalently the Lamé parameters $\lambda$ and $\mu$), external loading specifications $f$ (e.g., body force and its direction), and boundary conditions (e.g., fixed region $\Gamma_{D}$). Details on the physical parameterization and setup are provided in the supplementary material.
III-B Implicit Feature Learning
For a point cloud dataset $\mathcal{D}=\left\{{{P}_{i}}\right\}_{i=1}^{\left|\mathcal{D}\right|}$, we use encoder ${\Phi}_{\theta}$} and decoder ${\Psi}_{\eta}$} to learn the implicit features of the point cloud $P_{i}$}. The encoder ${\Phi}_{\theta}$} takes the $P_{i}$} as input and maps it to an $m$-dimensional latent vecoter $z_{\eta}$}, represented as ${{\Phi}_{\theta}}:{{\mathbb{R}}^{n\times 3}}\to{{\mathbb{R}}^{m}},m\ll n$, where we use the network backbones of PointNet [5] or DGCNN [7] as ${\Phi}_{\theta}$} to learn the point cloud features, respectively. Subsequently, we concatenate latent vector $z_{\eta}$} with an arbitrary query point $q$ in 3D space to obtain feature vector $z_{\eta}^{q}={z_{\eta}}\oplus q$. Query points are randomly sampled within the normalized object bounding box, with an additional subset sampled near the surface to better capture geometric details. Then, the decoder ${\Psi}_{\eta}$} takes $z_{\eta}^{q}$ as input and output an implicit function $\mathbb{S}={{\Psi}_{\eta}}\left(q,{{\Phi}_{\theta}}\left({{P}_{i}}\right)\right)=\left\{{\delta_{i}}|{\delta_{i}}\left({z_{\eta}^{q}}\right)\in\mathbb{R},\forall q\in{{\mathbb{R}}^{3}}\right\}$. ${\delta_{i}}\left({z_{\eta}^{q}}\right)$ represents the shortest distance from the query point $q$ to the point cloud $P_{i}$}. The decoder ${\Psi}_{\eta}$} consists of 3-layer of MLP and outputs a 1-dimensional feature vector.
@@ III-B Implicit Feature Learning
III-C Physical Information Awareness
@@
For a point cloud ${P}_{i}$}, we first use 3D Delaunay [53] to convert the ${P}_{i}$} into an object composed of multiple tetrahedra. It is an extension of Delaunay triangulation (which generates triangular meshes in 2D) to 3D space, producing a mesh composed of tetrahedral elements. Compared to hexahedral meshes, tetrahedral meshes better fit free-form objects and support complex topological structures, providing the volumetric elements required for physical modeling, such as deformation gradients and stress tensor computation. Meanwhile, directly using point coordinates lacks structural information, making it difficult to establish the strain–stress relationships necessary for constructing the subsequent physics-informed loss ${\mathcal{L}}_{pi}$. The advantages of using tetrahedral structures over point-based representations will be further validated in the ablation studies. The meshing process is shown in Figure 3.
@@
Figure 3: The point cloud is first converted into a coarse mesh form through Delaunay triangulation. Then, we removed the larger tetrahedral meshes, further refining them into the final tetrahedral representation.
@@
Subsequently, the tetrahedral mesh is processed by the mesh processor $\mathrm{T}$, which encodes the local geometric and topological structure of each tetrahedron into a unified format suitable for input to the encoder ${\Phi}_{\theta}$}. The mesh processor $\mathrm{T}$ consists of 3-layer of convolutional networks, bridging the mesh data and encoder ${\Phi}_{\theta}$}. The encoder ${\Phi}_{\theta}$} maps the processed data to the $m$-dimensional latent space to obtain the latent vector $z_{phys}$}. $z_{phys}$} is concatenated with the deformation parameters $\tau$ as the input $z_{phy}^{\tau}$ to decoder ${\Psi}_{phys}$}, enabling the model to learn local-to-whole deformation patterns under varying physical conditions during training. The parameters in $\tau$ have clear physical meanings and characterize typical elastic responses under small deformations. Considering variations in point-cloud density and deformation scale, we further apply an adaptive scaling to $\tau$ based on the size of the generated tetrahedral mesh and the initial deformation magnitude, reducing manual tuning and improving robustness and generalization.The decoder ${\Psi}_{phys}$} is implemented as a five-layer MLP and outputs an $n\times 3$ displacement field, where $n$ denotes the number of input points.
@@
Based on the configuration of deformation parameters $\tau$, we obtain the ground-truth displacement field $u$ using the FEM. Finally, based on the displacement field $\hat{u}$ predicted by the network and the ground-truth displacement field $u$ obtained through computation, we design two loss functions: the data-fidelity loss and the physics-informed loss. The data-fidelity loss guides the network to predict the displacement of each point in the deformed point cloud, explicitly modeling the relationship between local deformations and the whole geometric shape, thereby revealing the influence of local features on the whole structure. The physics-informed loss derives the external force distribution from the predicted local deformation data and compares it with the true external force distribution, ensuring that the network’s output satisfies the physical law of force equilibrium. Below, we provide a detailed description of these two loss functions.
Data-fidelity Loss. Its objective is to guide the network in accurately predicting the displacement field $\hat{u}$ deformed tetrahedral mesh object and comparing it with the ground truth displacement field u{u}, enabling it to precisely predict local geometric changes. For a tetrahedral mesh with undeformed nodal positions $x$, the deformed configuration is determined by node-wise displacements, i.e., $x^{\prime}=x+u$. Since local geometric variations (e.g., stretching, compression, and bending patterns) arise from these nodal motions, estimating u provides an effective way to preserve geometric consistency of deformation at the element level. From the perspective of mechanics, forces are determined by deformation, and deformation is parameterized by the displacement field $u$. In the general dynamic setting, the deformation trajectory $x\left(t\right)$ can be obtained by extremizing the action functional:
@@
where $S(x)$ denotes the action functional of the system over the time interval $\left[{0,T}\right]$. Here, $\frac{1}{2}{{{\dot{x}}}^{T}}M\dot{x}$ denotes the kinetic energy, with $M$ being the finite element mass matrix, $E(x)$ represents the potential energy (elastic deformation potential energy), and ${{x}^{T}}\left({{f}_{e}}+{{f}_{d}}\right)$ corresponds to the external force action. ${f}_{e}$} are external forces and ${f}_{d}$} are dissipative frictional forces.
The above presents the general variational formulation for dynamic problems. To analyze deformation under static linear elasticity over the spatial domain $\Omega$}, we typically solve for the displacement field $u$ by minimizing the total potential energy:
@@
where $\Omega$} represents the volume region of the object. $S(u)$ is the total potential energy as a functional of the displacement field. $\frac{1}{2}\sigma\colon\varepsilon(u)$ represents the energy stored in the tetrahedral mesh due to elastic deformation. After the mesh is subjected to force, the resulting energy is calculated by the stress $\sigma$ and strain $\varepsilon$. $f\cdot u$ represents the potential energy associated with external body forces under the displacement field, which is the work done by the external force under the displacement field $u$. Finally, we design the following data-fidelity loss:
@@
To ensure that the predicted deformation adheres to physical laws rather than merely fitting geometric targets, we introduce a physics-informed constraint governed by the fundamental force equilibrium equation [54]. In static elasticity, for any volume element within the mesh to remain in a physically stable state, the divergence of its internal stress tensor $\sigma$ must balance the external body forces $f$. This governing equation is formally defined as:
$\nabla\cdot\sigma+f=0\quad.$ (6)
Given the network predicted displacement field $\hat{u}$, we compute the stress field $\sigma$ and minimize the equilibrium residual $\left\|{\nabla\cdot\sigma+f}\right\|$ over all tetrahedra. This force-balance constraint couples neighboring elements and encourages globally consistent displacement predictions.
Specifically, we reconstruct the stress field $\sigma$ from the predicted vertex displacements using standard kinematics and Hooke’s law [55]. First, each tetrahedron consists of four vertices with initial coordinates ${{x}_{1}},{{x}_{2}},{{x}_{3}},{{x}_{4}}$ and deformed coordinates ${{{x}^{\prime}}_{1}},{{{x}^{\prime}}_{2}},{{{x}^{\prime}}_{3}},{{{x}^{\prime}}_{4}}$ (obtained from the network predictions). We construct the initial shape matrix ${X}=\left[{{x_{2}}-{x_{1}}\quad{\rm{}}{x_{3}}-{x_{1}}\quad{\rm{}}{x_{4}}-{x_{1}}}\right]$ and the deformed shape matrix ${X^{\prime}}=\left[{{{x^{\prime}}_{2}}-{{x^{\prime}}_{1}}\quad{\rm{}}{{x^{\prime}}_{3}}-{{x^{\prime}}_{1}}\quad{\rm{}}{{x^{\prime}}_{4}}-{{x^{\prime}}_{1}}}\right]$. Then, we construct the deformation gradient $F$ based on the shape matrices $X$ and $X^{\prime}$:
@@
where ${F}^{T}$ denotes the transpose of the deformation gradien and $I$ is the identity matrix. Under linear elastic materials, the constitutive relation (Hooke’s law) yields the 3D stress tensor as:
@@
where $\lambda$ and $\mu$ is the Lamé constant. $Tr\left(\varepsilon\right)$ is the trace of the strain tensor, representing the volumetric deformation. Finally, we define the physics-informed loss by averaging the $l_{2}$} norm of the equilibrium residual over all tetrahedra: